# A Catalog of First-Order Electroweak Phase Transitions in the Standard Model Effective Field Theory

### v1

@@ 1 Introduction
There have been a number of previous studies of the electroweak phase transition in the SMEFT or in more restricted models with dimension-six operators. The first such study was in Ref. [14], with later work in, e.g., Refs. [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25] mainly focusing on tree-level barriers arising from the dimension-six operator, which allows a negative Higgs quartic coupling $\lambda$. This, however, requires a low cutoff scale: it has been questioned whether such a low cutoff is reasonable [26, 24], and it has been argued dimension eight operators also become necessary [21].
The last few years have seen a rapid development in the methods used to study first-order phase transitions, allowing us to consistently describe models in which a barrier can arise in different ways. The modern methods are based on EFT principles, which gives an intuitive physical picture and a built-in region of applicability [27]—each form of barrier considered here comes with its own perturbative description. There is ongoing interest and active research in refining perturbation theory for describing first-order phase transitions [23, 28, 29, 30, 31, 32, 33, 34, 35], often in comparison with lattice studies [36, 37, 38, 39, 40, 41].
@@ 2 Theoretical framework
$V_{\mathrm{LO}}(\phi_{3})$ =?${=}\frac{1}{2}m_{3}^{2}\phi_{3}^{2}-\frac{1}{16\pi}g_{3}^{3}\phi_{3}^{3}+\frac{1}{4}\lambda_{3}\phi_{3}^{4}+\frac{1}{6}C^{\varphi}_{3}\phi_{3}^{6},$ (2.18)
@@ 3 The Catalog
We will take the gauge coupling $g$ of our gauge theory to be our power-counting parameter, and we will in this section study the simpler SU(2)+Higgs+$\phi^{6}$ theory as preparation for the full SMEFT in the next section. We do not include factors of $\pi$ here, as they are not important for deriving perturbative schemes. We will also assume that the corresponding 3D gauge coupling $g_{3}$ scales as $g_{3}\sim g\sqrt{T}$, which follows from the leading order matching relations [43]. Then by balancing the different terms in the potential, we can look for solutions among the different $n$’s that allow the different terms to be of sufficient size. There are also extra constraints on these powers coming from perturbativity. See appendix C for the details.
@@
[…]
@@
$6\displaystyle 3+3n_{\phi}<$ 6 $\implies n_{\phi}<1,$ (3.10)
$0\displaystyle 1+n_{\phi}>$ 0 $\implies n_{\phi}>-1,$ (3.11)
$0\displaystyle 3+n_{\phi}>$ 0 $\implies n_{\phi}>-3.$ (3.12)
@@
This exercise in power counting was relatively straightforward. With more possible coefficients in the potential, the counting can become more complicated. In the next subsection, we catalog the result of balancing the possible terms in $\mathrm{SU}(2)$+Higgs$+\phi^{6}$. More details can be found in appendix C.
@@ 3.1 Overview of the catalog
SC RB $g^{3}$ $g^{\frac{7}{2}}$ $g^{\frac{1}{2}}$ $<g^{\frac{3}{2}}$
SC RSB $g^{4}$ $g^{3}$ $g^{0}$ $<g^{3}$
$m_{3}\ll M\sim\pi T$ RSB $g^{4}$ $g^{2}$ $g^{-1}$ $<g^{6}$
@@
Figure 1: An overview of how the 3D parameter space can be sliced into different power-counting regions. Each region matches an entry in the catalog, in table 2. The white stars denote power-countings for which the SMEFT is estimated to predict the correct Higgs mass and vev, see subsection 4.3. The black dot indicating the SM is based on lattice studies [42] together with the DR matching relations. On the axes, note the “logarithmic” scale, and the hatch marks which emphasize that the axes have been cut. Finally, note that these regions are a projection which might lose some information, as the values of $m_{3}^{2}$ are not included, and that certain supercooled countings have been excluded as they overlap with other regions. As such, the figure should only be taken as a guiding sketch.
@@ 4 Mapping the parameter space of the SMEFT
Note that we now return to the 4D theory—it is important to note that we constructed the catalog in table 2 purely from the point of view of the 3D EFT that describes the scalar field. It is meant to primarily be an exploration of the possible kinds of barriers in gauge theories with a $\phi^{6}$ term. This means that we cannot guarantee that each counting can actually be realized by matching to the parameters of the corresponding 4D QFT, though we will explore this question in this subsection. We will now include the full SMEFT electroweak sector with the SU(2)$\times$×U(1) gauge group and corresponding gauge couplings $g$ and $g^{\prime}$, and in addition the Yukawa coupling $y_{t}$ of the top quark. These interactions are included in the parameter scans that will follow, and will appear in the matching conditions of the 3D and 4D theories.
@@ 4.1 The setup
We do not include dimension-8 operators in our analysis. Consider adding a term $-C^{8}\phi^{8}$ to the potential (4.6). Using Eqs. (4.74.8) it is then straightforward to see that if the Wilson coefficient $C^{8}$ takes its natural, large, value $\Lambda^{-4}$, it will be suppressed by one factor of $\lambda$ compared to the other terms in the potential. It will therefore not play a role in the formation of the barrier or in determining the value of the minimum.
@@ 4.2 Viable parameter space
[…]
@@
Figure 2: The points found by our scan, with better likelihood points on top. The points are colored according to whether they fall within the $1\sigma$, $2\sigma$, or $3\sigma$ confidence regions compared to the SM null hypothesis. The black line is the simplified prediction for which parameters get the correct Higgs mass and vev, based on the analysis in (A.15). The shaded regions reflect which power countings from table 2 will be relevant for studying the phase transition in those regions, according to the analysis in sections 4.3 and 4.4.
@@ 4.3 Mapping 4D to 3D
Except for the final two relations in eqs. 4.26 and 4.27, which were derived by hand, these relations were derived using the Mathematica code DRalgo [67], and cross-checked with those in [23]. The formula for $m_{3}^{2}$ is not the same as the one we used for $m^{2}_{\mathrm{eff}}$ in [10]: we here include the $C^{\varphi}$ contribution, and neglect the $C^{\varphi D}$ and $C^{\varphi\Box}$ contributions—a reflection of the power counting discussion after eqs. 4.7, 4.8, 4.9 and 4.10. Inspecting the relations in eqs. 4.20, 4.21, 4.22, 4.23, 4.24, 4.25, 4.26 and 4.27 tells us a few things. First, $g_{3}$ and $g^{\prime}_{3}$ are proportional to their 4D counterparts. This allows us to estimate their sizes just from the sizes of the corresponding 4D quantities: $g_{3}/\sqrt{T}\sim g$.
The coefficients $m_{3}^{2}$ and $\lambda_{3}$ are somewhat more difficult, because it is possible for cancellations to take place that disconnect the sizes of the 3D parameters from the 4D parameters. We have already seen an example of this in section 3.1 where we noted that $m_{3}^{2}$ can be smaller than $g^{2}T^{2}$ due to cancellations between $-m^{2}$ and the high-temperature corrections $\sim g^{2}T^{2}$. In principle the same could happen with $\lambda_{3}$, a possibility we explore in appendix B. It turns out this possibility is not perturbatively viable, and would require fine-tuning of the parameters. We will hence ignore this scenario and simply estimate $\lambda_{3}/T\sim\lambda$.
@@
Because this parameter point is expected to support a tree-level barrier, and hence $\lambda\sim g^{2}$, we emphasize that it is only CW-like: it does not feature radiative symmetry breaking. We can think of it as “effectively” scale-invariant: the only scale present comes from the cutoff-scale that defines the Wilson coefficient $C^{\varphi}$.
@@ 4.4 Identifying first-order phase transitions
$\lambda\sim g^{4}:$ $\quad 0\leq\lambda\leq 0.008$ (4.36c)
@@ 5.1 Electroweak baryogenesis and sphaleron transitions
Sphaleron transitions [71]—fluctuations in the electroweak fields which can violate $B-L$ and hence change baryon number—play a key role in electroweak baryogenesis. First, such transitions help generate a nonzero baryon number in the plasma just outside the bubble walls. Second, the sphaleron transitions must be sufficiently suppressed in the broken phase so that they do not wash out the generated baryon number.
@@
$-\log\Gamma_{\mathrm{sph}}\simeq\frac{E_{\mathrm{sph}}}{T}\sim\frac{v_{c}}{gT}\sim\frac{v_{3}}{g_{3}},$ (5.2)
@@
Hierarchy Shorthand $\alpha/\alpha_{\mathrm{RB}}$ $\beta/H\Big{/}(\beta/H)_{\mathrm{RB}}$ $-\log\Gamma_{\mathrm{sph}}$
$m_{3}\sim M\ll\pi T$ TLB $1$ $g^{\frac{3}{2}}$ $g^{-1}$
SC TLB (1) $1$ $g^{\frac{3}{2}}$ $g^{-1}$
@@
SC SQTLB (2) $1$ $g^{\frac{1}{4}}$ $g^{-\frac{5}{4}}$
RB $1$ $1$ $g^{-1}$
RB 6 $1$ $1$ $g^{-1}$
@@ 5.2 Gravitational wave signal
The phase transition strength $\alpha$ plays a key role in determining the energy budget of GWs in the plasma. Many measures of this strength are discussed in [82], including using the pressure difference, the trace anomaly, or the pseudotrace anomaly [87, 88]—with the pseudotrace anomaly being the most accurate choice, though a bit more complicated to calculate. Luckily for our case, the difference between the trace anomaly and the pseudotrace anomaly arises beyond leading order [89], and hence we can use the trace anomaly to estimate the size of $\alpha$. The definition of $\alpha$ is then
$\alpha=\frac{1}{p_{\mathrm{sym}}^{\prime}}\left(\Delta V_{3}-\frac{1}{4}\frac{\mathrm{d}\Delta V_{3}}{\mathrm{d}\log T}\right),$ (5.4)
where $p_{\mathrm{sym}}\sim T^{4}$ is the pressure in the symmetric phase, and $\Delta$ denotes the difference between the symmetric and broken phases: $\Delta X=X_{\mathrm{sym}}-X_{\mathrm{bro}}$.
@@
However, by inspecting the “blob” in figure 2 of [69], we conclude that the blob is quite large. And because our power-counting estimate is rather crude, if a given power-counting in fact moves this blob in one of the two promising directions, it might be enough for the blob to somewhat overlap the sensitivity region. With this widened condition, also the TLB power-counting scheme is interesting.
@@ 6 Discussion
This paper has laid the necessary groundwork for some future studies. Lattice studies can be performed, to better understand the phase portrait of $\mathrm{SU}(2)+\phi^{6}+$Higgs. Such studies would improve our understanding of the parameter regions of the SMEFT that describe first-order phase transitions. Furthermore, the various perturbative prescriptions outlined in this paper can be pushed to higher orders, as has been done for the radiative barrier case [41]. This would improve our understanding of the convergence of these prescriptions, and would increase the robustness of the GW signal calculations against their inherent large uncertainties—refining the accuracy of our leading order estimates. The catalog could also be used as a stepping-off point to study the electroweak phase transition in models with a $\phi^{6}$ term and new fields.
@@ Acknowledgments
We thank Andreas Ekstedt, Oliver Gould, and Tuomas Tenkanen for enlightening discussions. Part of this work was inspired from discussions held at the workshop “Thermal Field Theory and Early Universe Phenomenology at the Botanical Garden” held at the Swedish Collegium for Advanced Study in June 2023. JL’s work was funded by E och K.G. Lennanders stipendiestiftelse. The computations were performed on the Google Cloud Platform under the research credits program. The figures were made with the Julia package Makie [92].

### v2

@@ 1 Introduction
There have been a number of previous studies of the electroweak phase transition in the SMEFT or in more restricted models with dimension-six operators. The first such study was in Ref. [14], with later work in, e.g., Refs. [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27] mainly focusing on tree-level barriers arising from the dimension-six operator, which allows a negative Higgs quartic coupling $\lambda$. This, however, requires a low cutoff scale: it has been questioned whether such a low cutoff is reasonable [28, 24], and it has been argued dimension eight operators also become necessary [21].
The last few years have seen a rapid development in the methods used to study first-order phase transitions, allowing us to consistently describe models in which a barrier can arise in different ways. The modern methods are based on EFT principles, which gives an intuitive physical picture and a built-in region of applicability [29]—each form of barrier considered here comes with its own perturbative description. There is ongoing interest and active research in refining perturbation theory for describing first-order phase transitions [23, 30, 31, 32, 33, 34, 35, 36, 37], often in comparison with lattice studies [38, 39, 40, 41, 42, 43, 44].
@@ 2 Theoretical framework
$V_{\mathrm{LO}}(\phi_{3})$ $\overset{?}{=}\frac{1}{2}m_{3}^{2}\phi_{3}^{2}-\frac{1}{16\pi}g_{3}^{3}\phi_{3}^{3}+\frac{1}{4}\lambda_{3}\phi_{3}^{4}+\frac{1}{6}C^{\varphi}_{3}\phi_{3}^{6},$ (2.18)
@@ 3 The Catalog
[…]
@@
$V(\phi_{3})$ $\gg g^{6}T^{3},\qquad$ $\text{Perturbativity},$ (3.7)
$g_{3}^{2}\phi_{3}^{2}$ $\ll T^{2},\qquad$ $\text{High-T expansion for gauge boson},$ (3.8)
$m_{3}^{2}$ $\ll T^{2},\qquad$ $\text{High-T expansion for scalar}.$ (3.9)
@@
$3+3n_{\phi}<6$ $\implies n_{\phi}<1,$ (3.10)
$1+n_{\phi}>0$ $\implies n_{\phi}>-1,$ (3.11)
$3+n_{\phi}>0$ $\implies n_{\phi}>-3.$ (3.12)
@@
[…]
@@ 3.1 Overview of the catalog
SC RB $g^{3}$ $g^{\frac{7}{2}}$ $g^{\frac{1}{2}}$ $g^{\frac{3}{2}}$
SC RSB $g^{4}$ $g^{3}$ $g^{0}$ $g^{3}$
$m_{3}\ll M\sim\pi T$ RSB $g^{4}$ $g^{2}$ $g^{-1}$ $g^{6}$
@@
Figure 1: An overview of how the 3D parameter space can be sliced into different power-counting regions. Each region matches an entry in the catalog, in table 2. The white stars denote power-countings for which the SMEFT is estimated to predict the correct Higgs mass and vev, see subsection 4.3. The black dot indicating the SM is based on lattice studies [45] together with the DR matching relations. On the axes, note the ”logarithmic” scale, and the hatch marks which emphasize that the axes have been cut. Finally, note that these regions are a projection which might lose some information, as the values of $m_{3}^{2}$ are not included, and that certain supercooled countings have been excluded as they overlap with other regions. As such, the figure should only be taken as a guiding sketch.
@@ 4 Mapping the parameter space of the SMEFT
[…]
@@ 4.1 The setup
We do not include dimension-8 operators in our analysis. Consider adding a term $C^{8}\phi^{8}$ to the potential (4.6). Using Eqs. (4.74.8) it is then straightforward to see that if the Wilson coefficient $C^{8}$ takes its natural, large, value $\Lambda^{-4}$, it will be suppressed by one factor of $\lambda$ compared to the other terms in the potential. It will therefore not play a role in the formation of the barrier or in determining the value of the minimum.
@@ 4.2 Viable parameter space
In other words, we mostly use the ($M_{Z}$, $\alpha$, $G_{F}$) input scheme, with one technical exception: We find $\lambda$ to ensure that the calculated Higgs mass matches the observed value of $125.09\,\text{GeV}$ within a tolerance of $1\,\text{GeV}$. With this approach then $\lambda$ takes values that are different from the SM in order to match the Higgs mass and be consistent with the input value of $v$.
@@
Figure 2: The points found by our scan, with better likelihood points on top. The points are colored according to whether they fall within the $1\sigma$, $2\sigma$, or $3\sigma$ confidence regions compared to the SM null hypothesis. The black line is the simplified prediction for which parameters get the correct Higgs mass and vev, based on the analysis in (A.15). The shaded regions reflect which power countings from table 2 will be relevant for studying the phase transition in those regions, according to the analysis in sections 4.3 and $4.$ The boundaries demarcating these regions are leading-order estimates based on the analysis in Sec. 4.4; they are expected to shift and theoretical uncertainties upon the inclusion of higher-order corrections.
@@ 4.3 Mapping 4D to 3D
Except for the final two relations in eqs. 4.26 and 4.27, which were derived by hand, these relations were derived using the Mathematica code DRalgo [70], and cross-checked with those in [23]. These matching relations can be extended beyond leading order, and in fact such exactness is necessary when performing high precision calculations. As we are here only interested in deriving power counting relations, we are content with not going beyond the leading order expressions. Any higher order corrections, being suppressed relative to the leading order, will not affect the power counting rules.
The formula for $m_{3}^{2}$ is not the same as the one we used for $m^{2}_{\mathrm{eff}}$ in [10]: we here include the $C^{\varphi}$ contribution, and neglect the $C^{\varphi D}$ and $C^{\varphi\Box}$ contributions—a reflection of the power counting discussion after eqs. 4.7, 4.8, 4.9 and 4.10. Inspecting the relations in eqs. 4.20, 4.21, 4.22, 4.23, 4.24, 4.25, 4.26 and 4.27 tells us a few things. First, $g_{3}$ and $g^{\prime}_{3}$ are proportional to their 4D counterparts. This allows us to estimate their sizes just from the sizes of the corresponding 4D quantities: $g_{3}/\sqrt{T}\sim g$.
The coefficients $m_{3}^{2}$ and $\lambda_{3}$ are somewhat more difficult, because it is possible for cancellations to take place that disconnect the sizes of the 3D parameters from the 4D parameters. We have already seen an example of this in section 3.1 where we noted that $m_{3}^{2}$ can be smaller than $g^{2}T^{2}$ due to cancellations between $m^{2}$ and the high-temperature corrections $\sim g^{2}T^{2}$. In principle the same could happen with $\lambda_{3}$, a possibility we explore in appendix B. It turns out this possibility is not perturbatively viable, and would require fine-tuning of the parameters. We will hence ignore this scenario and simply estimate $\lambda_{3}/T\sim\lambda$.
@@
Because this parameter point is expected to support a tree-level barrier, and hence $\lambda\sim g^{2}$, we emphasize that it is only CW-like: it does not feature radiative symmetry breaking. We can think of it as ”effectively” scale-invariant: the only scale present comes from the cutoff-scale that defines the Wilson coefficient $C^{\varphi}$.
@@ 4.4 Identifying first-order phase transitions
$\lambda\sim g^{4}:$ $\quad 0\leq\lambda\leq 0.008.$ (4.36c)
@@ 5.1 Electroweak baryogenesis and sphaleron transitions
Sphaleron transitions [74]—fluctuations in the electroweak fields which can violate BLB-L and hence change baryon number—play a key role in electroweak baryogenesis. First, such transitions help generate a nonzero baryon number in the plasma just outside the bubble walls. Second, the sphaleron transitions must be sufficiently suppressed in the broken phase so that they do not wash out the generated baryon number.
@@
$\log\Gamma_{\mathrm{sph}}\simeq\frac{E_{\mathrm{sph}}}{T}\sim\frac{v_{c}}{gT}\sim\frac{v_{3}}{g_{3}},$ (5.2)
@@
Hierarchy Shorthand $\alpha/\alpha_{\mathrm{RB}}$ $\beta/H\Big/(\beta/H)_{\mathrm{RB}}$ $\log\Gamma_{\mathrm{sph}}$
$m_{3}\sim M\ll\pi T$ TLB 11 $g^{\frac{3}{2}}$ $g^{-1}$
SC TLB (1) 11 $g^{\frac{3}{2}}$ $g^{-1}$
@@
SC SQTLB (2) 11 $g^{\frac{1}{4}}$ $g^{-\frac{5}{4}}$
RB 11 11 $g^{-1}$
RB 6 11 11 $g^{-1}$
@@ 5.2 Gravitational wave signal
The phase transition strength $\alpha$ plays a key role in determining the energy budget of GWs in the plasma. Many measures of this strength are discussed in [85], including using the pressure difference, the trace anomaly, or the pseudotrace anomaly [90, 91]—with the pseudotrace anomaly being the most accurate choice, though a bit more complicated to calculate. Luckily for our case, the difference between the trace anomaly and the pseudotrace anomaly arises beyond leading order [92], and hence we can use the trace anomaly to estimate the size of $\alpha$. The definition of $\alpha$ is then the trace anomaly divided by $3w$, where $w$ is the enthalpy density, which for a bag model equation of state satisfies $w=(4/3)\rho$, where $\rho=T\partial p/\partial T$ is the energy density and $p$ is the pressure. We then have (using $V_{4}=TV_{3}$)
$\alpha=\left(\frac{\partial p_{\mathrm{sym}}}{\partial T}\right)^{-1}\left(\Delta V_{3}-\frac{1}{4}\frac{\mathrm{d}\Delta V_{3}}{\mathrm{d}\log T}\right),$ (5.4)
where $p_{\mathrm{sym}}\sim T^{4}$ is the pressure in the symmetric phase and $\Delta$ denotes the difference between the symmetric and broken phases: $\Delta X=X_{\mathrm{sym}}-X_{\mathrm{bro}}$.
@@
However, by inspecting the ”blob” in figure 2 of [72], we conclude that the blob is quite large. And because our power-counting estimate is rather crude, if a given power-counting in fact moves this blob in one of the two promising directions, it might be enough for the blob to somewhat overlap the sensitivity region. With this widened condition, also the TLB power-counting scheme is interesting.
@@
5.3 Higgs pair production
Let us finally briefly comment on implications for Higgs pair production at the LHC. This process is intimately connected to the physics of the electroweak phase transition, as it involves the Higgs quartic coupling present in the Higgs potential. In the SM there are two one-loop diagrams that contribute at leading order, one involving both the Higgs-top quark coupling and the triple Higgs coupling $\lambda_{HHH}$, which is in turn proportional to the quartic coupling parameter $\lambda$, the other involving two powers of the Higgs-top coupling. There is destructive interference between the two diagrams and therefore a rather small cross-section. In the SMEFT, both diagrams receive corrections from dimension-six operators, and in addition there are several new diagrams involving SMEFT operators both at tree level and one-loop level. In particular the triple Higgs coupling in the SMEFT involves the quartic $\lambda$ as well as the three Wilson coefficients $C^{\varphi}$, $C^{\varphi\Box}$ and $C^{\varphi D}$. As we have seen, for scenarios with a first-order phase transition, $\lambda$ can have values substantially different from the SM value, as well as the opposite sign from the SM. Moreover, the triple Higgs vertex in the SMEFT is momentum-dependent, proportional to $C^{\varphi\Box}$ and $C^{\varphi D}$. Precise measurements of the differential cross-section and its momentum dependence at the high-luminosity LHC could therefore test the scenarios presented here. A detailed study of Higgs pair production is beyond the scope of this paper, and we refer the reader to our previous paper [10] for more details, and, e.g., [93] for further discussion. [^9] We note, however, that most sources discussing Higgs pair production in the SMEFT choose an input scheme that fixes the Higgs quartic coupling to the same value as in the SM, making the analysis of the potential more complicated.
@@ 6 Discussion
It is important to comment on the expected impact of NLO corrections on our findings. As we note in Secs. 4.4 and 5.2, our analysis represents a leading-order framework, and we have emphasized that higher-order corrections are necessary for precision predictions. We anticipate that NLO corrections will refine, rather than invalidate, the qualitative classification presented in our catalog. Their primary effect would be to quantitatively shift the boundaries between the phenomenological regions depicted in figure 2, i.e. the sharp vertical lines, which represent our LO estimates based on the analysis in Sec. 4.4, would be replaced by theoretical uncertainty bands. However, the existence of these distinct scenarios, i.e. tree-level barriers, radiative barriers, is a direct consequence of the features of the potential which we identify using our power-counting framework. Our work provides a map of the SMEFT landscape, which is a necessary prerequisite for targeted NLO studies in the future.
For example, lattice studies can be performed, to better understand the phase portrait of $\mathrm{SU}(2)+\phi^{6}+$Higgs. Such studies would improve our understanding of the parameter regions of the SMEFT that describe first-order phase transitions. Furthermore, the various perturbative prescriptions outlined in this paper can be pushed to higher orders, as has been done for the radiative barrier case [43]. This would improve our understanding of the convergence of these prescriptions, and would increase the robustness of the GW signal calculations against their inherent large uncertainties—refining the accuracy of our leading order estimates. The catalog could also be used as a stepping-off point to study the electroweak phase transition in models with a $\phi^{6}$ term and new fields.
@@ Acknowledgments
We thank Andreas Ekstedt, Oliver Gould, and Tuomas Tenkanen for enlightening discussions. Part of this work was inspired from discussions held at the workshop ”Thermal Field Theory and Early Universe Phenomenology at the Botanical Garden” held at the Swedish Collegium for Advanced Study in June 2023. JL’s work was funded by E och K.G. Lennanders stipendiestiftelse. The computations were performed on the Google Cloud Platform under the research credits program. The figures were made with the Julia package Makie [96].